### v1

# GeoWarp: An automatically differentiable and GPU-accelerated implicit MPM framework for geomechanics based on NVIDIA Warp

Abstract
The material point method (MPM), a hybrid Lagrangian–Eulerian particle method, is increasingly used to simulate large-deformation and history-dependent behavior of geomaterials. While explicit time integration dominates current MPM implementations due to its algorithmic simplicity, such schemes are unsuitable for quasi-static and long-term processes typical in geomechanics. Implicit MPM formulations are free of these limitations but remain less adopted, largely due to the difficulty of computing the Jacobian matrix required for Newton-type solvers, especially when consistent tangent operators should be derived for complex constitutive models. In this paper, we introduce GeoWarp—an implicit MPM framework for geomechanics built on NVIDIA Warp—that exploits GPU parallelism and reverse-mode automatic differentiation to compute Jacobians without manual derivation. To enhance efficiency, we develop a sparse Jacobian construction algorithm that leverages the localized particle–grid interactions intrinsic to MPM. The framework is verified through forward and inverse examples in large-deformation elastoplasticity and coupled poromechanics. Results demonstrate that GeoWarp provides a robust, scalable, and extensible platform for differentiable implicit MPM simulation in computational geomechanics.
keywords:
Material point method , Implicit method , Automatic differentiation , Differentiable simulation , Large deformation , Geomechanics
journal:
1 Introduction
The material point method (MPM) [sulsky1994particle] is a continuum particle method that combines Lagrangian and Eulerian descriptions. It tracks the state of materials using a set of particles in a Lagrangian framework, while solving the governing equations on a background grid following an Eulerian approach. This hybrid Lagrangian–Eulerian formulation allows MPM to simulate large-deformation kinematics of history-dependent materials without suffering from mesh distortion. This capability makes MPM particularly well suited for modeling geomaterials (e.g. soils, rocks, and snow) that undergo large deformations. As a result, it has been increasingly adopted in computational geomechanics to model the behavior of geomaterials and their interactions with various objects and processes (e.g. [zabala2011progressive, yerro2015material, kularathna2021semi, dunatunga2015continuum, gaume2018dynamic, li2021three, seyedan2021solid, dunatunga2022modelling, jiang2022hybrid, jiang2023erosion, liang2023revealing, zhao2023coupled, yu2024semi, zhao2024mapped]).
Most existing MPM implementations employ explicit time integration schemes—particularly the explicit Euler method—due to their algorithmic simplicity, ease of implementation, and suitability for dynamic problems. Accordingly, explicit MPM has been widely adopted in open-source codes such as Uintah-MPM [germain2000uintah], NairnMPM [nairn2015], CB-Geo MPM [kumar2019scalable], Karamelo [de2021karamelo], fMPM-solver [wyser2020fast], GeoTaich [shi2024geotaichi], Anura3D [Anura3D2025], MaterialPointSolver.jl [huo2025high], and Matter [blatny2025matter]. Despite their advantages, explicit schemes are only conditionally stable and impose severe limitations on the time step size. Furthermore, explicit methods do not allow control over numerical error per time step. Due to these reasons, explicit MPM is undesirable for quasi-static or long-term problems prevalent in geomechanics [lacroix2020life].
Implicit time integration schemes, in contrast, are unconditionally stable and permit error control at each time step. These features make them more suitable for quasi-static or long-term processes, and implicit solvers have been widely employed in geomechanics (e.g. [young2009liquefaction, borja2010continuum, choo2015stabilized, choo2016hydromechanical, choo2018cracking]). However, implementing implicit MPM is considerably more challenging, as it requires solving a nonlinear system of equations at every time step. This is typically achieved using Newton’s method, which in turn demands accurate derivation and implementation of the Jacobian matrix to ensure convergence. Deriving the Jacobian matrix is particularly tedious and error-prone when the formulation involves complex constitutive models, as the consistent tangent operator—defined as the derivative of stress with respect to strain consistent with the time integration algorithm—should also be computed [simo1985consistent]. In the context of MPM, only a limited number of studies have pursued implicit formulations (e.g. [guilkey2003implicit, beuth2011solution, wang2016development, charlton2017igimp, coombs2020ample, coombs2020lagrangian, zhao2020stabilized, chandra2024mixed]). Among these, AMPLE [coombs2020ample] provides an open-source implementation of implicit MPM with clear documentation and example applications. However, as noted by the authors, AMPLE is primarily intended for conceptual demonstration and is not designed for high-performance computing or large-scale simulation. To date, no open-source, high-performance implementation of implicit MPM tailored specifically for computational geomechanics has been made available.
Automatic differentiation (AD) offers a promising pathway toward implementing implicit solvers and enabling differentiable simulation. By constructing a computational graph during the forward pass and applying the chain rule in reverse during the backward pass, AD allows for the exact and automatic computation of gradients. This capability has significantly impacted simulation-based fields, including computational mechanics, where AD is increasingly used to simplify the implementation of nonlinear solvers. In the context of the finite element method (FEM), recent studies have demonstrated that AD can greatly reduce development effort by eliminating the need to manually derive and implement Jacobian matrices for complex, path-dependent constitutive models (e.g. [vigliotti2021automatic, xue2023jax, dummer2024robust]). Beyond implementation efficiency, AD also enables gradient-based optimization workflows, including inverse analysis, control, and learning. In the MPM community, several recent frameworks have successfully integrated AD with explicit MPM solvers. For example, ChainQueen [hu2019chainqueen], DiffTaichi [hu2019difftaichi], and PlasticineLab [huang2021plasticinelab] demonstrate differentiable simulations for soft robotics and control applications by embedding AD into GPU-accelerated MPM engines. However, these frameworks are primarily built for dynamic simulations using explicit time integration. The extension of AD to implicit MPM remains largely unexplored, despite its potential to simplify the implementation of Newton-type solvers and support inverse modeling in geomechanical problems.
Leveraging recent advances in AD, here we introduce GeoWarp—an open-source, high-performance framework for implicit MPM tailored to computational geomechanics built on NVIDIA Warp [warp2022]. Warp is a Python-based simulation platform that supports AD and just-in-time compilation for efficient execution on both CPUs and GPUs. Its differentiable kernel-based programming model enables automatic and accurate gradient computation, while its Python interface facilitates rapid development. These features make Warp a suitable foundation for constructing differentiable solvers for large-scale mechanics problems.
At the core of GeoWarp is a fully implicit MPM solver integrated with AD. This integration eliminates the need for manually deriving and implementing Jacobian matrices, which are required for Newton-type solvers. In geomechanics, where constitutive models are often nonlinear and history-dependent, consistent tangent operators can be particularly difficult to derive and implement correctly. By recording the full computational graph of the simulation, GeoWarp enables automatic Jacobian computation, substantially lowering the barrier to implementing robust implicit MPM methods.
To address the computational cost typically associated with AD, we introduce a Jacobian construction strategy that leverages the sparsity pattern inherent in MPM. By exploiting the locality of particle–grid interactions, the method reduces reverse-mode evaluations to a small, fixed number of backward passes independent of problem size. This approach enables efficient large-scale simulation on modern GPUs, even for three-dimensional problems. Beyond forward simulation, the automatic differentiability of GeoWarp facilitates inverse modeling tasks such as material parameter identification and integration with gradient-based optimization and learning-based methods.
The framework is verified through forward and inverse examples in large-deformation elastoplasticity and coupled poromechanics. Results demonstrate that GeoWarp provides a robust, scalable, and extensible platform for differentiable implicit MPM simulations in computational geomechanics. The implementation is released as an open-source codebase to support reproducibility and future development: https://github.com/choo-group/GeoWarp.
2 Material point method formulation
This section outlines the implicit MPM formulation adopted in this work. We begin by stating the initial–boundary value problem governing large-deformation mechanics, followed by its spatial discretization using the material point method. The resulting nonlinear system is then solved using a fully implicit scheme with Newton’s method. For brevity, we focus on describing the essential components of MPM, referring readers to comprehensive references for further details [jiang2016material, zhang2016material, fern2019material, nguyen2023material].
2.1 Problem statement
Consider a deformable body occupying a domain $\Omega$ with boundary $\partial\Omega$. The boundary is partitioned into a Dirichlet portion $\partial_{\bm{u}}\Omega$ where displacements are prescribed, and a Neumann portion $\partial_{\bm{t}}\Omega$ where tractions are applied. These subsets satisfy the standard conditions $\partial_{\bm{u}}\Omega\cap\partial_{\bm{t}}\Omega=\emptyset$ and $\overline{\partial_{\bm{u}}\Omega\cup\partial_{\bm{t}}\Omega}=\partial\Omega$. The problem is defined over a time interval $\mathcal{T}:=(0,T]$, with $T>0$.
To accommodate large deformations, we adopt a finite deformation framework that distinguishes between the reference and current configurations. A material point is identified by its position vector in the reference configuration, $\bm{X}$, and its position vector in the current configuration, $\bm{x}$. The displacement vector is then defined as $\bm{u}:=\bm{x}-\bm{X}$. The deformation gradient is defined as
$\bm{F}:=\dfrac{\partial\bm{x}}{\partial\bm{X}}=\bm{1}+\dfrac{\partial\bm{u}}{\partial\bm{X}},$ (1)
where $\bm{1}$ denotes the second-order identity tensor. The Jacobian,
$J:=\det(\bm{F}),$ (2)
represents the ratio of the current differential volume, $\mathrm{d}v$, to the reference differential volume, $\mathrm{d}V$.
In the current configuration, the momentum balance under quasi-static conditions is given by
$\operatorname{\nabla\cdot}\bm{\sigma}(\bm{F})+\rho\bm{g}=\bm{0}\quad\text{in}\quad\Omega\times\mathcal{T},$ (3)
[…]
$\bm{u}=\hat{\bm{u}}\quad$ $\text{on}\quad\partial\Omega_{\bm{u}}\times\mathcal{T},$ (4)
$\bm{n}\cdot\bm{\sigma}=\hat{\bm{t}}\quad$ $\text{on}\quad\partial\Omega_{\bm{t}}\times\mathcal{T},$ (5)
with $\hat{\bm{u}}$ and $\hat{\bm{t}}$ denoting the prescribed boundary displacement and traction, respectively. The initial condition is given by
$\bm{u}=\bm{u}_{0}\quad\text{at}\quad t=0,$ (6)
where $\bm{u}_{0}$ denotes the initial displacement field. To close the formulation, a constitutive law relating the Cauchy stress tensor $\bm{\sigma}$ to the deformation gradient $\bm{F}$ should be specified. In this study, we consider several constitutive models commonly used in geomechanics. For brevity, we omit their detailed formulations and refer the reader to standard references on constitutive models (e.g. [de2011computational, borja2013plasticity]).
2.2 Implicit MPM discretization
For the MPM discretization of the problem, we represent the domain as a collection of particles (material points) and introduce a background computational grid that interacts with the particles. The particles carry state variables such as stress and volume, following a Lagrangian description. In contrast, the governing equations are discretized and solved on the background grid, which follows an Eulerian description. To couple these two descriptions, field quantities are transferred between the particles and grid nodes. Let us use the subscripts $(\cdot)p$ and $(\cdot)i$ to denote quantities associated with particle $p$ and grid node $i$, respectively. The mapping of an arbitrary variable $f$ from particles to nodes can be written as
$f_{i}=\sum_{p=1}^{n_{p}}w_{i,p}m_{p}f_{p}/M_{i}\,,\quad M_{i}:=\sum_{p=1}^{n_{p}}w_{i,p}m_{p}\,.$ (7)
where $n_{p}$ is the number of material points influencing node $i$, $m_{p}$ is the mass of particle $p$, and $w_{i,p}$ is the weighting function associated with node $i$ evaluated at the position of particle $p$. These weighting functions serve the same role as shape functions in the finite element method.
For the MPM weighting functions, the original MPM formulation [sulsky1994particle] employs linear finite element shape functions; however, such functions are prone to interpolation errors when particles cross cell boundaries. To alleviate this issue, several enhanced interpolation schemes have been developed, including the Generalized Interpolation Material Point (GIMP) [bardenhagen2004generalized] method, Convected Particle Domain Interpolation (CPDI) [sadeghirad2011convected, sadeghirad2013second], B-spline-based approaches [steffen2008examination, steffen2008analysis], and Moving Least Squares (MLS) interpolation [hu2018moving]. In this work, we adopt the contiguous GIMP (cpGIMP) scheme [wallstedt2008evaluation], which has been widely used in conjunction with implicit MPM formulations [charlton2017igimp, coombs2020ample, coombs2020lagrangian]. The one-dimensional weighting function and its gradient follow the formulation presented by Coombs and Augarde [coombs2020ample]. In multiple dimensions, the weighting functions and their gradients are constructed as tensor products of the corresponding one-dimensional components along each spatial direction.
At each time step, the MPM update follows a four-stage procedure, as illustrated in Fig. 1 and described below:
  1. 1.
    Particle-to-grid (P2G) transfer: Particle quantities are mapped to the background grid using the selected weighting functions.
  2. 2.
    Nodal update: The governing equations are solved on the grid to compute nodal displacements (or velocities, depending on the formulation).
  3. 3.
    Grid-to-particle (G2P) transfer: Updated nodal values are interpolated back to the particles to recover particle-level kinematics.
  4. 4.
    Particle update: Particle positions and state variables are updated accordingly. The background grid is then reset in preparation for the next time step.
Figure 1: MPM update procedure.
Once the above procedure is completed, the simulation advances to the next time step, repeating the four-stage process until the end of the analysis.
In the second step, the governing equation is solved in its variational form. Applying the standard procedure to the momentum balance gives
$-\int_{\Omega}\operatorname{\nabla^{s}}\bm{\eta}:\bm{\sigma}\,\mathrm{d}v+\int_{\Omega}\bm{\eta}\cdot\rho\bm{g}\,\mathrm{d}v+\int_{\partial\Omega_{\bm{t}}}\bm{\eta}\cdot\hat{\bm{t}}\,\mathrm{d}a=0,$ (8)
where $\bm{\eta}$ denotes the variation of the displacement field, $\operatorname{\nabla^{s}}$ denotes the symmetric gradient operator evaluated in the current configuration, and $\mathrm{d}a$ denotes the differential surface area in the current configuration. The variational equation can be discretized using the Galerkin method. The resulting discrete form is evaluated as
$\operatorname*{\bm{\mathsf{A}}}_{p}\left(\bm{f}_{{\rm int},p}(\bm{u}_{p})\right)=\operatorname*{\bm{\mathsf{A}}}_{p}\left(\bm{f}_{{\rm ext},p}\right),$ (9)
where $\operatorname*{\bm{\mathsf{A}}}_{p}$ denotes the global assembly over all particles, and $\bm{f}_{{\rm int},p}$ and $\bm{f}_{{\rm ext},p}$ are the internal and external force vectors, respectively, defined as
$\bm{f}_{{\rm int},p}$ $:=\left(\operatorname{\nabla^{s}}\bm{w}_{p}\right)^{\top}:\bm{\sigma}_{p}V_{p},$ (10)
$\bm{f}_{{\rm ext},p}$ $:=\bm{w}_{p}^{\top}\rho_{p}\bm{g}V_{p}+\bm{w}_{p}^{\top}\hat{\bm{t}}_{p}V_{p},$ (11)
with $\bm{w}_{p}$ and $\operatorname{\nabla^{s}}\bm{w}_{p}$ denoting the weighting functions and their symmetric gradients evaluated at grid nodes influenced by particle $p$, and $V_{p}$ denoting the particle volume.
In this work, we solve the discrete governing equation using the implicit (backward) Euler method. Let superscripts $(\cdot)^{n}$ and $(\cdot)^{n+1}$ denote quantities at time steps $n$ and $n+1$, respectively. Given all relevant quantities at step $n$, our goal is to compute the displacement field $\bm{u}^{n+1}$ at step $n+1$. Since the governing equation is generally nonlinear in $\bm{u}^{n+1}$, we make use of Newton’s method to solve the residual equation
$\bm{r}(\bm{u}^{n+1})=\operatorname*{\bm{\mathsf{A}}}_{p}\left(\bm{f}_{{\rm int},p}(\bm{u}_{p}^{n+1})-\bm{f}_{{\rm ext},p}^{n+1}\right)\rightarrow\bm{0},$ (12)
where $\bm{r}(\bm{u}^{n+1})$ is the global residual vector assembled from particle contributions. At each iteration $k$, we compute the update by solving the linear system
$-\bm{J}^{n+1,k}\Delta\bm{u}^{n+1,\,k}=\bm{r}(\bm{u}^{n+1,k}),$ (13)
where the Jacobian matrix is defined as
$\bm{J}^{n+1,k}:=\dfrac{\partial\bm{r}(\bm{u}^{n+1,k})}{\partial\bm{u}^{n+1,k}}.$ (14)
The iterations continue until the relative residual norm satisfies the prescribed convergence criterion:
$\dfrac{\|\bm{r}(\bm{u}^{n+1,k})\|}{\|\bm{r}(\bm{u}^{n+1,0})\|}\leq\text{tol},$ (15)
where tol denotes the tolerance. Once convergence is achieved, the displacement field is updated using the increment $\Delta\bm{u}^{n+1,k}$.
The convergence behavior of Newton’s method is highly sensitive to the accuracy of the Jacobian matrix. In computational mechanics, the primary difficulty in evaluating the Jacobian lies in computing the consistent tangent operator, defined as the derivative of the incremental stress with respect to the incremental strain (or an equivalent measure) [simo1985consistent]. For relatively simple constitutive models, this operator can be derived analytically. However, for complex models commonly used in geomechanics, the derivation and implementation of consistent tangents is often tedious, error-prone, and model-specific. To overcome this limitation, we leverage AD to compute Jacobian matrices in a manner that is accurate, general, and fully automated. The details of our AD-based approach are presented in the following section.
Remark 1.
In addition to the single-phase mechanics formulation described above, we consider a coupled poromechanics formulation based on the $\bm{u}$$p$ approach, which is widely used in geomechanics to model fluid flow and solid deformation in saturated porous media (e.g. [white2008stabilized, choo2018enriched, choo2018large, choo2019stabilized]). This formulation introduces the pore pressure $p$ as an additional primary variable, augmenting the solid momentum balance with a fluid mass conservation equation. The resulting system captures two-way coupling between fluid diffusion and solid deformation. The $\bm{u}$$p$ formulation is also implemented in GeoWarp based on the MPM formulation described in Zhao and Choo [zhao2020stabilized]. The reader is referred to that work for full details of the governing equations and numerical implementation.